A farmer is going to divide her 60 acre farm between two crops. Seed for crop A costs $20 per acre. Seed for crop B costs $40 per acre. The farmer can spend at most $1400 on seed.
If crop B brings in a profit of $270 per acre, and crop A brings in a profit of $120 per acre, how many acres of each crop should the farmer plant to maximize her profit?

Answer by ikleyn(52783)

.
A farmer is going to divide her 60 acre farm between two crops.
Seed for crop A costs $20 per acre. Seed for crop B costs $40 per acre. The farmer can spend at most $1400 on seed.
If crop B brings in a profit of $270 per acre, and crop A brings in a profit of $120 per acre,
how many acres of each crop should the farmer plant to maximize her profit?
~~~~~~~~~~~~~~~~~

Seed for crop B is twice as expensive as seed for crop A, but gives the profit which is more than twice as big

as the profit of crop A (compare  270/120 = 2.25  with  40/20 = 2 !).


    +---------------------------------------------------------------+
    |   THEREFORE, it is clear, that the most aggressive strategy   |
    |   is to sow as much area with crop B as possible.             |
    +---------------------------------------------------------------+


The possible area to sow with crop B, within the budget, is  1400/40 = 35 acres.


So, the most profitable solution is to sow 35 acres with crop B and do not sow crop A, at all.


The expected profit is then  35*270 = 9450 dollars, and it is maximal possible profit at given conditions.

Solved.